OpenStudy (anonymous):

Solve for x. 17 = –13 – 8x

OpenStudy (anonymous):

17+13 = -8x 30 = -8x -30/8 = x x = 15/4

OpenStudy (anonymous):

sorry x = -15/4

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw